# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 7 Lattice spin systems with discrete gauge symmetries
In this section we review the phase diagram and critical behavior of gauge spin systems with discrete gauge group, focusing mainly on lattice ${\mathbb{Z}}_{2}$-gauge $N$-vector models, obtained by minimally coupling $N$-component real fields with ${\mathbb{Z}}_{2}$-gauge fields [231]. As discussed in Sec. 3.3 they are paradigmatic models showing LGW× phase transitions. They display different phases, characterized by the spontaneous breaking of the global O($N$) symmetry and by the different topological properties of the ${\mathbb{Z}}_{2}$-gauge correlations, see, e.g., Refs. [41, 20]. They are relevant in several different contexts, see, e.g., Refs. [302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314]. In particular, they are relevant for transitions in nematic liquid crystal [308, 310, 311, 315] and for systems with fractionalized quantum numbers, see, e.g., Refs. [313, 304].
@@ 7.1 The ${\mathbb{Z}}_{2}$-gauge $N$-vector models
The ${\mathbb{Z}}_{2}$-gauge $N$-vector model is a lattice $N$-vector model with local ${\mathbb{Z}}_{2}$ gauge invariance. From the point of view of the symmetries, it can be interpreted as an $N$-vector model, which is symmetric under O($N$) = ${\mathbb{Z}}_{2}\otimes\mathrm{SO}(N)$ transformations, in which the ${\mathbb{Z}}_{2}$ symmetry is gauged, giving rise to a ${\mathbb{Z}}_{2}$-gauge theory coupled with scalar vector fields. Its lattice Hamiltonian is reported in Eq. (17). It is invariant under global SO($N$) transformations ${\bm{s}}_{\bm{x}}\to V{\bm{s}}_{\bm{x}}$ with $V\in\mathrm{SO}(N)$, and local ${\mathbb{Z}}_{2}$ gauge transformations,
@@
Due to the ${\mathbb{Z}}_{2}$ gauge invariance, the correlation function of the vector variables ${\bm{s}}_{x}$, $G_{s}({\bm{x}},{\bm{y}})=\langle{\bm{s}}_{\bm{x}}\cdot{\bm{s}}_{\bm{y}}\rangle$, trivially vanishes for ${\bm{x}}\neq{\bm{y}}$. For $N\geq 2$, the spontaneous breaking of the global SO($N$) symmetry is signaled by the condensation of the gauge-invariant bilinear operator $R_{\bm{x}}^{ab}$,
@@ 7.2 Phase diagram and critical behavior of ${\mathbb{Z}}_{2}$-gauge $N$-vector models with $N\geq 2$
The phase diagram for $N\geq 2$ is sketched in Fig. 12, see, e.g., Refs. [41, 20, 231]. Two spin-disordered phases are present for small $J$: a small-$K$ phase, in which both spin and ${\mathbb{Z}}_{2}$-gauge variables are disordered (DD), and a large-$K$ phase in which the ${\mathbb{Z}}_{2}$-gauge variables order (DO). For large $J$ there is a single ordered phase in which both spins and gauge variables order (O). These phases are separated by three transition lines, which meet at one point, see Fig. 12.292929An estimate of the intersection point $(K_{\star},J_{\star})$ of the three lines for small values of $N$ ($N\lesssim 5$ say) is reported in Ref. [231]: $K_{\star}\approx 0.761-0.003N\approx 0.75$, $J_{\star}\approx 0.23$. Note that in the ordered phase, once a representative of the minimum energy configuration is chosen, i. e. ${\bm{s}}_{\bm{x}}={\bm{s}}_{{\bm{x}}+\hat{\mu}}$ and $w_{{\bm{x}},\mu}=1$ (up to gauge transformations and in the thermodynamic limit), there is no residual gauge symmetry, hence we do not expect the presence of topological phases or transitions, cf. Sec. 3.5.3.
@@ 7.2.2 RPN-1-like transitions along the DD-O transition line
The order parameter of the DD-O transitions is the bilinear operator $R^{ab}_{\bm{x}}$ defined in Eq. (70). For $N=2$, if the transition is continuous, it behaves as an XY vector field, so its RG dimension $Y_{R}$ coincides with the RG dimension $Y_{V,{\rm XY}}=(1+\eta_{\rm XY})/2$ of the vector field in the standard XY model. Thus, at continuous transitions along the DD-O line for $N=2$, we have $Y_{R}=Y_{V,{\rm XY}}=0.519088(22)$ (using the estimates of $\eta_{\rm XY}$ reported in B). The transitions along the DD-O line for $N=2$ have been also investigated numerically [231], confirming that continuous transitions behave as predicted by the LGW approach, but also observing that the continuous transitions turn into first-order ones for $K=K_{\rm tri}$, with $K_{\rm tri}<K^{*}$, i.e., before the intersection point (at $K=K^{\star}$) of the transition lines. Some numerical results are reported in the left panel of Fig. 13.
@@ 7.2.3 O($N$)× transitions along the DO-O transition line
As discussed in Sec. 7.1, the breaking of the SO($N$) symmetry along the DO-O line is signaled by the condensation of the operator $R^{ab}_{\bm{x}}$ defined in Eq. (70). Along the DO-O line this operator, as well as all gauge-invariant operators, have the same critical behavior as in the conventional $N$-vector model without gauge invariance [231]. The equivalence of the gauge-invariant correlations in O($N$) models and along the DO-O line implies that the gauge modes do not drive the critical behavior for finite large values of $K$. Moreover, these transitions have the same effective description as the conventional O($N$) ones, although there is no appropriate gauge-invariant order parameter. These transitions are therefore O($N$)× transitions, according to the classification reported in Sec. 3.1. Note that these transitions differ from the RPN-1-like transitions that occur along the DD-O line, which are also LGW transitions but with a different, gauge-invariant, order parameter, see Sec. 7.2.2.
@@
In O($N$)× transitions, the main issue is the identification of the non-gauge-invariant order parameter. In noncompact models, local order parameters are identified by introducing an appropriate gauge fixing, the Lorenz gauge, see Sec. 5 and, particular, Sec. 5.1.2. In the context of compact models, one can exploit a stochastic gauge-fixing procedure [232, 231], as discussed below in Sec. 7.3. The introduction of this peculiar gauge fixing allows one to identify the spin variable ${\bm{s}}_{\bm{x}}$ as the order parameter of the transition, as in standard $N$-vector systems. Numerical results for $N=1,2,3$ reported in Refs. [232, 231] fully confirm these scenarios. In the gauge-fixed theory, along the DO-O line the vector field ${\bm{s}}_{\bm{x}}$ behaves as in the standard $N$-vector model.
Let us finally focus on the ${\mathbb{Z}}_{2}$-gauge XY ($N=2$) model, which is the only case in which continuous transitions may occur along all three transition lines. It is interesting to note that the continuous transitions along the DD-O and DO-O lines are all expected to belong to the 3D XY universality class. However, this does not imply that the critical behavior is the same, as the two transition lines have different effective LGW descriptions. This is clearly demonstrated by the different critical behavior of the operator $R^{ab}_{\bm{x}}$ defined in Eq. (70). As discussed in Sec. 7.2.2, the bilinear operator $R_{\bm{x}}^{ab}$ is the effective XY order parameter at continuous DD-O transitions, so its RG dimensions is the same as that of the vector field in the XY model: $Y_{R}=Y_{V,{\rm XY}}=(1+\eta_{\rm XY})/2\approx 0.519$. On the other hand, along the DO-O line, gauge-invariant quantities behave as in the O($N$) vector model. Therefore, the operator $R_{\bm{x}}^{ab}$ behaves as a spin-two tensor operator at a standard XY transition. It follows that its RG dimension equals the RG dimension $Y_{T,{\rm XY}}$ of the spin-2 operator in the XY model (numerical estimates of this quantity are reported in Refs. [321, 250, 173, 322]): $Y_{R}=Y_{T,{\rm XY}}=1.23629(11)$. These predictions have been verified numerically in Ref. [231]. As an example, some results are reported Fig. 13, which confirm the identification of the DO-O transitions as O($N$)× continuous transitions with a gauge-dependent vector order parameter, and of the DD-O as LGW transitions with a gauge-invariant spin-two order parameter.
@@ 7.3 The stochastic gauge fixing
To define a local O($N$)× order parameter, a conceptually different procedure from the usual gauge-fixing strategies was proposed in Refs. [232, 231], introducing a stochastic gauge fixing that breaks gauge invariance without changing gauge-invariant correlation functions. This procedure generalizes the variational approach that is used in the context of lattice gauge theories and is usually called Landau gauge fixing. [^33] In a lattice gauge theory with gauge fields $U_{{\bm{x}},\mu}$, the Landau gauge is defined as the set of gauge fields that maximize $\sum_{{\bm{x}}\mu}\hbox{Re Tr }U_{{\bm{x}},\mu}$, see, e.g., Ref. [323]. This approach is numerically difficult to apply and, in general, it suffers from the so-called Gribov problem [324, 325]: there may be several degenerate or quasidegenerate maxima. In the continuum limit, it goes over to the gauge-fixing condition $\partial_{\mu}A_{\mu}=0$, that we have named Lorenz gauge, but it is also referred to as Landau gauge. In continuum theories Landau gauge and Lorenz gauge are often used interchangeably, although the name “Landau gauge” is often preferred for Euclidean field theories, “Lorenz gauge” for theories defined in Minkowski space. To be consistent with the definitions of Sec. 5, we would call it Lorenz gauge. We review it below, in the context of the ${\mathbb{Z}}_{2}$-gauge $N$-vector model.
@@
where $\hat{\bm{s}}_{\bm{x}}$ and $\hat{\sigma}_{{\bm{x}},\mu}$ correspond to the fields obtained by performing a gauge transformation with gauge function $w_{\bm{x}}$, cf. Eq. (69). The average $[A({\bm{s}}_{\bm{x}},\sigma_{{\bm{x}},\mu})]$ is then averaged over the fields ${\bm{s}}_{\bm{x}}$ and $\sigma_{{\bm{x}},\mu}$ using the original Hamiltonian (17), i.e.
@@
One can easily see that gauge-invariant observables are invariant under this gauge-fixing procedure. In this approach, one defines the vector correlation function as
@@
The resulting model with the added variables $w_{\bm{x}}$ is a quenched random-bond Ising model [151, 326] with a particular choice of bond distribution. Quenched random-bond Ising models have various phases—disordered, ferromagnetic, and glassy phases—depending on the temperature (whose role is played here by $1/\gamma$), the amount of randomness of the bond distribution, and its spatial correlations, see, e.g., Refs. [151, 326, 327, 328, 329, 330, 331, 332, 333, 334, 335, 336, 337, 338, 339, 340]. In particular, the present model is expected to undergo a quenched transition for $\gamma=\gamma_{c}(J,K)$. The transition separates a disordered phase for $\gamma<\gamma_{c}(J,K)$ from a large-$\gamma$ phase, which a priori can be ferromagnetic or glassy. If $J$ and $K$ belong to the DO-O transition line, the large-$\gamma$ phase turns out to be ferromagnetic. [^36] Note that the two-point function of $w_{\bm{x}}$ vanishes for non-coincident points, due to the generalized gauge invariance, Eq. (77). The nature of the large-$\gamma$ phase can be determined by considering correlation functions of the so-called overlap $O_{\bm{x}}=w_{\bm{x}}^{(1)}w_{\bm{x}}^{(2)}$, where $w_{\bm{x}}^{(1)}$ and $w_{\bm{x}}^{(2)}$ represent two different configurations sampled with the same probability $e^{-H_{w}}/Z_{w}$, i.e. corresponding to the same values of $\sigma_{{\bm{x}},\mu}$ and $s_{\bm{x}}$. Along the DO-O line, for $\gamma>\gamma_{c}(J,K)$ the Binder cumulant of the overlap variable approaches $1$, and the overlap susceptibility diverges linearly with the volume, indicating that the large-$\gamma$ phase is a standard ferromagnetic phase. Thus, the long-distance behavior of the variables $w_{\bm{x}}$ is the same for all $\gamma>\gamma_{c}(J,K)$: The variables $w_{\bm{x}}$ simply make uncorrelated short-range fluctuations around the minimum configurations obtained for $\gamma\to\infty$. It is thus natural to conjecture that $\gamma$ is an irrelevant parameter, i.e., that the critical behavior of the gauge-fixed quantities is the same for any $\gamma>\gamma_{c}(J,K)$ along the DO-O transition line. In RG language, $1/\gamma$ represents an irrelevant perturbation of the $\gamma=\infty$ fixed point. The irrelevance of $\gamma$ is conjectured to be a general feature of the stochastic gauge fixing, which holds for any $N$ (the numerical analyses of Refs. [232, 231] confirm this conjecture). It is important to stress the significant advantage of the stochastic gauge fixing with respect to traditional gauge-fixing approaches. As we can work at fixed $\gamma$, there is no need to determine the minima or maxima of some gauge-fixing function, bypassing the problem of the Gribov copies [324, 325].
@@
Figure 14: (Adapted from Ref. [232]) FSS analyses (see A.3) of MC data for the $\mathbb{Z}_{2}$-gauge $N$-vector model with $N=3$, using the stochastic gauge fixing with parameter $\gamma=0.3$. Simulations at fixed $K=1$ across the DO-O transition line. (Left) Plot of the vector Binder parameter $U_{V}$ defined in terms of ${\bm{s}}_{\bm{x}}$, versus $\xi_{V}/L$, where $\xi_{V}$ is the second-moment correlation length computed by using the vector correlation function $G_{V}$. Data collapse on the $N$-vector ($N=3$) universal curve (see [271] for a parametrization), showing that the vector degrees of freedom behaves as at a standard O(3) transition. (Right) Plot of the vector susceptibility $\chi_{V}$ determined using $G_{V}$ versus the ratio $\xi_{V}/L$. The susceptibility is rescaled according to Eq. (153), using the exponent $\eta_{H}=0.03784$ of the Heisenberg O(3) universality class (see B for a list of estimates of $\eta_{H}$). The nice collapse of the data supports the O(3) nature of the vector modes in the presence of a stochastic gauge fixing.

### v3

@@ 7 Lattice spin systems with discrete gauge symmetries
In this section we review results on lattice gauge spin systems with discrete gauge group, focusing mainly on lattice ${\mathbb{Z}}_{2}$-gauge $N$-vector models, obtained by minimally coupling $N$-component real fields with ${\mathbb{Z}}_{2}$-gauge fields [262]. As discussed in Sec. 3.3 they are paradigmatic models undergoing LGW× phase transitions.
Lattice ${\mathbb{Z}}_{2}$-gauge $N$-vector models are relevant in several different contexts, see, e.g., Refs. [330, 331, 332, 333, 334, 335, 336, 337, 338, 339, 340, 341, 342]. In particular, they are relevant for transitions in nematic liquid crystal [336, 338, 339, 343, 344, 345] and for systems with fractionalized quantum numbers, see, e.g., Refs. [341, 332].
We discuss their phase diagrams and critical behaviors, considering separately multicomponent ($N\geq 2$) and $N=1$ models. Multicomponent models display three different phases, characterized by the spontaneous breaking of the global O($N$) symmetry and by the different topological properties of the ${\mathbb{Z}}_{2}$-gauge correlations, see, e.g., Refs. [41, 20]. We also discuss the possibility of uncovering critical gauge-dependent correlations at thier LGW× transitions by implementing an appropriate stochastic gauge fixing. The particular case $N=1$ corresponds to the so-called ${\mathbb{Z}}_{2}$ Higgs model, which satisfies an exact duality relation [346]. Unlike multicomponent models, it presents only two phases separated by two Ising-like transition lines, which meet at an XY multicritical point.
The final part of this section presents a survey of results for other models with discrete gauge symmetries. In particular, we discuss whether there is an effective enlargement of the gauge symmetry at the critical point, as it occurs in the lattice ${\mathbb{Z}}_{Q}$ gauge models for a sufficiently large $Q$, see Sec. 3.5.2.
@@ 7.1 The ${\mathbb{Z}}_{2}$-gauge $N$-vector models
The ${\mathbb{Z}}_{2}$-gauge $N$-vector model is a lattice model with local ${\mathbb{Z}}_{2}$ gauge invariance. From the point of view of the symmetries, it can be interpreted as an $N$-vector model, which is symmetric under O($N$) = ${\mathbb{Z}}_{2}\otimes\mathrm{SO}(N)$ transformations, in which the ${\mathbb{Z}}_{2}$ symmetry is gauged, giving rise to a ${\mathbb{Z}}_{2}$-gauge theory coupled with scalar vector fields. Its lattice Hamiltonian is reported in Eq. (17). It is invariant under global SO($N$) transformations ${\bm{s}}_{\bm{x}}\to V{\bm{s}}_{\bm{x}}$ with $V\in\mathrm{SO}(N)$, and local ${\mathbb{Z}}_{2}$ gauge transformations,
@@
Due to the ${\mathbb{Z}}_{2}$ gauge invariance, the vector correlation function $G_{s}({\bm{x}},{\bm{y}})=\langle{\bm{s}}_{\bm{x}}\cdot{\bm{s}}_{\bm{y}}\rangle$, trivially vanishes for ${\bm{x}}\neq{\bm{y}}$. For $N\geq 2$, the spontaneous breaking of the global SO($N$) symmetry is signaled by the condensation of the gauge-invariant bilinear operator $R_{\bm{x}}^{ab}$,
@@ 7.2 Phase diagram and critical behavior of ${\mathbb{Z}}_{2}$-gauge $N$-vector models with $N\geq 2$
The phase diagram for $N\geq 2$ is sketched in Fig. 12, see, e.g., Refs. [41, 20, 262, 347, 348]. Two spin-disordered phases are present for small $J$: a small-$K$ phase, in which both spin and ${\mathbb{Z}}_{2}$-gauge variables are disordered (DD), and a large-$K$ phase in which the ${\mathbb{Z}}_{2}$-gauge variables order (DO). For large $J$ there is a single ordered phase in which both spins and gauge variables order (O). These phases are separated by three transition lines, which meet at one point, see Fig. 12.303030An estimate of the intersection point $(K_{\star},J_{\star})$ of the three lines for small values of $N$ ($N\lesssim 5$ say) is reported in Ref. [262]: $K_{\star}\approx 0.761-0.003N\approx 0.75$, $J_{\star}\approx 0.23$. Note that in the ordered phase, once a representative of the minimum energy configuration is chosen, i.e., ${\bm{s}}_{\bm{x}}={\bm{s}}_{{\bm{x}}+\hat{\mu}}$ and $w_{{\bm{x}},\mu}=1$ (up to gauge transformations and in the thermodynamic limit), there is no residual gauge symmetry. Therefore, we do not expect the presence of topological phases or transitions, see the discussion reported in Sec. 3.5.3.
@@ 7.2.2 RPN-1-like transitions along the DD-O transition line
The order parameter of the DD-O transitions is the bilinear operator $R^{ab}_{\bm{x}}$ defined in Eq. (73). For $N=2$, if the transition is continuous, it behaves as an XY vector field, so its RG dimension $Y_{R}$ coincides with the RG dimension $Y_{V,{\rm XY}}=(1+\eta_{\rm XY})/2$ of the vector field in the standard XY model. Thus, at continuous transitions along the DD-O line for $N=2$, we have $Y_{R}=Y_{V,{\rm XY}}=0.519088(22)$ (using the estimates of $\eta_{\rm XY}$ reported in B). The transitions along the DD-O line for $N=2$ have also been investigated numerically [262], confirming that continuous transitions behave as predicted by the LGW approach, but also observing that the continuous transitions turn into first-order ones for $K=K_{\rm tri}$, with $K_{\rm tri}<K^{*}$, i.e., before the intersection point (at $K=K^{\star}$) of the transition lines. Some numerical results are reported in the left panel of Fig. 13.
@@ 7.2.3 O($N$)× transitions along the DO-O transition line
As discussed in Sec. 7.1, the breaking of the SO($N$) symmetry along the DO-O line is signaled by the condensation of the operator $R^{ab}_{\bm{x}}$ defined in Eq. (73). Along the DO-O line this operator, as well as all gauge-invariant operators, have the same critical behavior as in the conventional $N$-vector model without gauge invariance [262]. The equivalence of the gauge-invariant correlations in O($N$) models and along the DO-O line implies that the gauge modes do not drive the critical behavior for finite large values of $K$. Moreover, these transitions have the same effective description as the conventional O($N$) ones, although there is no appropriate gauge-invariant order parameter. These transitions are therefore O($N$)× transitions, according to the classification reported in Sec. 3.1. Note that these transitions differ from the RPN-1 transitions that occur along the DD-O line, which are also LGW transitions but with a different, gauge-invariant, order parameter, see Sec. 7.2.2.
@@
In O($N$)× transitions, the main issue is the identification of the non-gauge-invariant order parameter. In noncompact models, local order parameters are identified by introducing an appropriate gauge fixing, the Lorenz gauge, see Sec. 5 and, particular, Sec. 5.1.2. In the context of compact models, one can exploit a stochastic gauge-fixing procedure [263, 262], as discussed below in Sec. 7.3. The introduction of this peculiar gauge fixing allows one to identify the spin variable ${\bm{s}}_{\bm{x}}$ as the order parameter of the transition, as in standard $N$-vector systems. Numerical results for $N=1,2,3$ reported in Refs. [263, 262] fully confirm this scenario. In the gauge-fixed theory, along the DO-O line the vector field ${\bm{s}}_{\bm{x}}$ behaves as in the standard $N$-vector model.
Let us finally focus on the ${\mathbb{Z}}_{2}$-gauge XY ($N=2$) model, which is the only case in which continuous transitions may occur along all three transition lines. It is interesting to note that the continuous transitions along the DD-O and DO-O lines are all expected to belong to the 3D XY universality class. However, this does not imply that the critical behavior is the same, as the two transition lines have different effective LGW descriptions. This is clearly demonstrated by the different critical behavior of the operator $R^{ab}_{\bm{x}}$ defined in Eq. (73). As discussed in Sec. 7.2.2, the bilinear operator $R_{\bm{x}}^{ab}$ is the effective XY order parameter at continuous DD-O transitions, so its RG dimensions is the same as that of the vector field in the XY model: $Y_{R}=Y_{V,{\rm XY}}=(1+\eta_{\rm XY})/2\approx 0.519$. On the other hand, along the DO-O line, gauge-invariant quantities behave as in the O($N$) vector model. Therefore, the operator $R_{\bm{x}}^{ab}$ behaves as a spin-two tensor operator at a standard XY transition. It follows that its RG dimension equals the RG dimension $Y_{T,{\rm XY}}$ of the spin-2 operator in the XY model (numerical estimates of this quantity are reported in Refs. [165, 157, 204, 353]): $Y_{R}=Y_{T,{\rm XY}}=1.23629(11)$. These predictions have been verified numerically in Ref. [262]. As an example, some results are reported Fig. 13, which confirm the identification of the DO-O transitions as O($N$)× continuous transitions with a gauge-dependent vector order parameter, and of the DD-O transitions as LGW transitions with a gauge-invariant spin-two order parameter.
@@ 7.3 The stochastic gauge fixing
To define a local O($N$)× order parameter, a novel procedure, which is conceptually different from the usual gauge-fixing strategies, was proposed in Refs. [263, 262]. It generalizes the variational approach that is used in the context of lattice gauge theories and is usually called Landau gauge fixing. [^34] In a lattice gauge theory with gauge fields $U_{{\bm{x}},\mu}$, the Landau gauge is defined as the set of gauge fields that maximize $\sum_{{\bm{x}}\mu}\hbox{Re Tr }U_{{\bm{x}},\mu}$, see, e.g., Ref. [354]. This approach is numerically difficult to apply and, in general, it suffers from the so-called Gribov problem [355, 356]: there may be several degenerate or quasidegenerate maxima. In the continuum limit, it goes over to the gauge-fixing condition $\partial_{\mu}A_{\mu}=0$, that we have named Lorenz gauge, but which is also referred to as Landau gauge. In continuum theories Landau gauge and Lorenz gauge are often used interchangeably, although the name “Landau gauge” is often preferred for Euclidean field theories, “Lorenz gauge” for theories defined in Minkowski space. To be consistent with the definitions of Sec. 5, we would call it Lorenz gauge. We review it below, in the context of the ${\mathbb{Z}}_{2}$-gauge $N$-vector model.
@@
where $\hat{\bm{s}}_{\bm{x}}$ and $\hat{\sigma}_{{\bm{x}},\mu}$ correspond to the fields obtained by performing a gauge transformation with gauge function $w_{\bm{x}}$ as in Eq. (72). The average $[A({\bm{s}}_{\bm{x}},\sigma_{{\bm{x}},\mu})]$ is then averaged over the fields ${\bm{s}}_{\bm{x}}$ and $\sigma_{{\bm{x}},\mu}$ using the original Hamiltonian (17), i.e.,
@@
One can easily see that gauge-invariant observables are invariant under this gauge-fixing procedure. In this approach, we can define the vector correlation function as
@@
The resulting model with the added variables $w_{\bm{x}}$ is a quenched random-bond Ising model [180, 358] with a particular choice of bond distribution. Quenched random-bond Ising models have various phases—disordered, ferromagnetic, and glassy phases—depending on the temperature (whose role is played here by $1/\gamma$), the amount of randomness of the bond distribution, and its spatial correlations, see, e.g., Refs. [180, 358, 359, 360, 361, 362, 363, 364, 365, 366, 367, 368, 369, 370, 371, 372, 373, 374]. In particular, the present model is expected to undergo a quenched transition at $\gamma=\gamma_{c}(J,K)$ for any $J$ and $K$. The transition separates a disordered phase for $\gamma<\gamma_{c}(J,K)$ from a large-$\gamma$ phase, which a priori can be ferromagnetic or glassy. If $J$ and $K$ belong to the DO-O transition line, the large-$\gamma$ phase turns out to be ferromagnetic.373737 Note that the two-point function of $w_{\bm{x}}$ vanishes for non-coincident points, due to the generalized gauge invariance, Eq. (80). The nature of the large-$\gamma$ phase can be determined by considering correlation functions of the so-called overlap $O_{\bm{x}}=w_{\bm{x}}^{(1)}w_{\bm{x}}^{(2)}$, where $w_{\bm{x}}^{(1)}$ and $w_{\bm{x}}^{(2)}$ represent two different configurations sampled with the same probability $e^{-H_{w}}/Z_{w}$, i.e. corresponding to the same values of $\sigma_{{\bm{x}},\mu}$ and $s_{\bm{x}}$. Along the DO-O line, for $\gamma>\gamma_{c}(J,K)$ the Binder cumulant of the overlap variable approaches $1$, and the overlap susceptibility diverges linearly with the volume, indicating that the large-$\gamma$ phase is a standard ferromagnetic phase. Thus, the long-distance behavior of the variables $w_{\bm{x}}$ is the same for all $\gamma>\gamma_{c}(J,K)$: The variables $w_{\bm{x}}$ simply make uncorrelated short-range fluctuations around the minimum configurations obtained for $\gamma\to\infty$. It is thus natural to conjecture that $\gamma$ is an irrelevant parameter, i.e., that the critical behavior of the gauge-fixed quantities is the same for any $\gamma>\gamma_{c}(J,K)$ along the DO-O transition line. In the RG language, $1/\gamma$ represents an irrelevant perturbation of the $\gamma=\infty$ fixed point. The irrelevance of $\gamma$ is conjectured to be a general feature of the stochastic gauge fixing, which holds for any $N$ (the numerical analyses of Refs. [263, 262] confirm this conjecture). It is important to stress the significant advantage of the stochastic gauge fixing with respect to traditional gauge-fixing approaches. As we can work at fixed $\gamma$, there is no need to determine the minima or maxima of some gauge-fixing function, bypassing the problem of the Gribov copies [355, 356].
@@
Figure 14: (Adapted from Ref. [263]) FSS analyses (see A.3 for a discussion of the expected scaling behavior) of MC data for the $\mathbb{Z}_{2}$-gauge $N$-vector model with $N=3$, using the stochastic gauge fixing with parameter $\gamma=0.3$. Simulations at fixed $K=1$ across the DO-O transition line. (Left) Plot of the vector Binder parameter $U_{V}$ defined in terms of ${\bm{s}}_{\bm{x}}$, versus $\xi_{V}/L$, where $\xi_{V}$ is the second-moment correlation length computed by using the vector correlation function $G_{V}$. Data collapse on the $N$-vector ($N=3$) universal curve (see Ref. [302] for a parametrization), showing that the vector degrees of freedom behave as at a standard O(3) transition. (Right) Plot of the vector susceptibility $\chi_{V}$ defined in terms of $G_{V}$ versus the ratio $\xi_{V}/L$. The susceptibility is rescaled according to Eq. (158), using the exponent $\eta_{H}=0.03784$ of the Heisenberg O(3) universality class (see B for a list of estimates of $\eta_{H}$). The nice collapse of the data supports the O(3) nature of the vector modes in the presence of a stochastic gauge fixing.